Marginal Cost Soni

The total cost for Soni corp. to manufacture q DVD players is given by the function C(x)= 6q^3 - q^2 + 7q + 90, where q represents the quantity of DVDs manufactured. Find the marginal cost for Soni Corp. when it manufactures 5 DVDs.
Answer #1

total cost c(x)=6q^3-q^2+7q+90

marginalcost dc/dq =18q^2-2q+7

marginal cost for soni corp when it manufactures 5 DVD's=18*5^2-2*5+7

=18*25 -3

=450-3

=447


answered by: mathboy